OpenStudy (anonymous):

What is the 7th term of the geometric sequence where a1 = 625 and a3 = 25? 0.0016 0.2 0.008 0.04

OpenStudy (lalaly):

\[a_n=a_1 \times r^{n-1}\]\[25=625 \times r^2\]\[r^2=0.04\]\[r=0.2\] \[a_7=625 \times 0.2^{7-1}\]

OpenStudy (anonymous):

So the answer would be 0.04?

OpenStudy (lalaly):

yep

OpenStudy (anonymous):

Thank you so much again!

OpenStudy (lalaly):

=)